Marcus has two sisters. The sum of the girls' ages is 13. The product of their ages is 36. How old is each sister?
Answer by checkley79(3341)

X+Y=13
X=13-Y
XY=36
(13-Y)Y=36
13Y-Y^2=36
-Y^2+13Y-36=0
Y^2-13Y+36=0
(Y-9)(Y-4)=0
Y-9=0
Y=9 YEARS OLD ANS. FOR ONE SISTER
X+9=13
X=13-9
X=4 YEARS OLD ANS. FOR THE OTHER SISTER
Y-4=0
Y=4 ANS. DITTO.
X+4=13
X=13-4
X=9 DITTO.
PROOF:
4+9=13
13=13
4*9=36
36=36